How do I solve the following using the substitution steps?
5x+2y=7
y=x-7
Thank you. TJ

Answer by richwmiller(17219)

This not a word problem.
The only words were your question.
plug (x-7) in for y in 5x+2y=7 to get
5x+2*(x-7)=7
5x+2x-14=7
7x=21
x=3
y=x-7
y=3-7
y=-4
(3,-4)